A mover drags a 10 kg trunk along a ramp inclined at 20° to the horizontal by applying a force of 200 N at an angle of 30° above the incline. If the force of frictional resistance acting on the trunk is 125 N, determine the acceleration of the trunk.
1
Expert's answer
2020-05-05T18:41:51-0400
"ma=F\cos{30}-mg \sin{20}-F_f"

"10a=200\cos{30}-10(9.8)\sin{20}-125"

"a=1.5\frac{m}{s^2}"